OpenStudy (anonymous):

You plan to build a house that is 1 1/2 times as long as it is wide. you want the land around the house to be 20 feet wider than the width of the house, and twice as long as the length of the house. 1) write an expression for the area of land surrounding the house. 2) if x=30, what is the area of the entire property?

OpenStudy (anonymous):

x = wide of house 3/2 x = length land width = x +20 land legth = 3x Area = 3x(x+20) = 3x^2 +60x